# Joint Graph Rewiring and Feature Denoising via Spectral Resonance

### v2

@@ A.5.3 Combining JDR and DIGL
We compare our method to DIGL (Gasteiger et al., 2019) in the main text. We use the personalized PageRank (PPR) diffusion kernel and the same top-$64$ values sparsening method as in JDR in all experiments. Figure 6 shows the additional results for DIGL on the synthetic datasets from the cSBM. Table 9 shows the results on the real-world homophilic datasets in the sparse splitting and Table 10 on the heterophilic datasets in the dense splitting. Here, in addition to the individual results for JDR and DIGL, the results for a combination of the two methods are also shown. For this purpose, the graph was first denoised with JDR and then diffused with DIGL. To do this, we fixed the hyperparameters of JDR and then tuned the parameter $\alpha$ of DIGL. We think this is interesting as both methods enhance the graph in different ways and thus should be combinable. In principle, this should also be possible for a combination of JDR with BORF or FoSR, but so far we have not been able to get this to work.
@@
Figure 6: Comparison of DIGL (Gasteiger et al., 2019) and JDR on the cSBM datasets in the sparse splitting. Results for (6(a)) GCN and (6(b)) GPRGNN as downstream models. The error bars indicate the $95\%$ confidence interval. As expected, DIGL is not really able to improve the performance of the GNNs in the heterophilic regime. It achieves the greatest improvement in the weak-graph regime and for strongly homophilic graphs, especially using GCN as downstream model. Another interesting observation is that for GCN and $\phi<0.25$ the curve of MLP corresponds exactly to the one of GCN +DIGL. The reason for this is that the hyperparameters found for DIGL ensure that the graph is ignored ($\alpha=1.0$), which means that the GCN then collapses to a simple MLP. For the more powerfull GPRGNN, on the other hand, DIGL is generally hardly able to improve performance, while JDR clearly increases the performance across all $\phi$.
A.5.4 Alignment
Here, we give a more detailed view on how much JDR actually increases alignment on cSBM and real-world datasets compared to the baseline methods. For cSBM, we can see in Figure 7 that DIGL only increases alignment in the homophilic regime. In the heterophilic regime it clearly decreases alignment. We expect this because it promotes connections among nodes at short diffusion distance. Also the random teleport probability found on these datasets is $1.0$, which results in a random uniformly connected graph. Similar to the results on the real-world datasets Cornell and Texas from the main text, we can see this in the classification performance in Figure 6. In the heterophilic regime, the performance of DIGL matches exactly the MLP, while in the homophilic regime, we can see some performance increases. With GPRGNN, DIGL is not really able to improve performance at all (except for $\phi=0$). In Figure 8, we can see that JDR increases alignment across all settings and more strongly than the baseline methods on the real-world graphs (except for Citeseer). DIGL also increases alignment on many homophilic graphs and on heteophilic Texas. We would like to note that when only rewiring a graph, increasing the alignment might not always be the best thing to do: If the graph is very good, it would be the features that should be made more aligned to the graph. But only JDR is able to do this, as it also denoises the features as well. It is also interesting that DIGL decreases the alignment for Cora and PubMed, but still achieves a good performance. This indicates that DIGL in this case improves the graph in a different way than JDR. So here, it should be possible to combine both methods to achieve even better performance. And indeed, A.5.3 shows that this is exactly the case for Cora, PubMed and Photo on GCN. FoSR and BORF do not visibly change the alignment, since their modifications to the graph are usually too small.
@@ A.5.4 Alignment
Figure 8: Alignment of the leading eigenspaces of graphs from homophilic (8(a)) and heterophilic (8(b)) real-world datasets. We compare the original graph (None) to the output of DIGL, BORF and JDR with the hyperparameters found on GCN. JDR increases the alignment in all settings and achieves the maximum alignment among rewiring algorithms in all settings except on the Citeseer dataset.
A.5.5 Spectral Clustering
@@
Figure 9: Separate results for using spectral clustering on a rewriting only $\bm{A}$ (9(a)) and denoising only $\bm{X}$ (9(b)) compared to full JDR. Note that for $\phi\in\{0.5,0.625,0.875\}$ we had to use additional graphs generated using cSBM with an average node degree of $d=10$ for spectral clustering of $\bm{A}$ to work in general and for $\phi=0.875$ also for JDR. The reason for this is that the graph is very sparse so it is not necessarily connected such that there is no guarantee that spectral clustering works. However a larger node degree does not improve the performance of spectral clustering in general, while it may for GNNs.
@@
Figure 10: Separate results for rewiring only $\bm{A}$ and denoising only $\bm{X}$ compared to full JDR. Results for (10(a)) GCN and (10(b)) GPRGNN. The error bars indicate the $95\%$ confidence interval.
@@ A.7 Hyperparameters
In this section we list all the hyperparameters used for the experiments to ensure the reproducibility of the results. They are also included in the code. In all experiments we use the Adam optimizer and the standard early stopping after 200 epochs from (Chien et al., 2021). Whenever we use a GCN, it uses two layers, a hidden dimension of $64$ and dropout with $0.5$. Whenever we use GPRGNN, we use a polynomial filter of order $10$ (corresponding to $10$ hops) and a hidden dimension of $64$. For JDR, we always keep the $64$ largest entries of the rewired adjacency matrix $\tilde{\bm{A}}$ per node. We justify this choice by the ablation in Figure 13.
cSBM. For synthetic data from the cSBM, we generally follow the hyperparameters from (Chien et al., 2021). GCN uses a learning rate of $0.01$ and weight decay with $\lambda=0.0005$. GPRGNN also uses a $\lambda=0.0005$ and both use ReLU non-linearity. On homophilic graphs ($\phi\geq 0$), GPRGNN uses a learning rate of $0.01$, a weight initialization $\alpha=0.1$ and dropout with $0.5$. For heterophilic graphs, it uses a learning rate of $0.05$, $\alpha=1.0$ and dropout $0.7$. The hyperparameters for JDR on the cSBM are shown in Table 17. We only tuned them using GCN as a downstream model, so for GPRGNN +JDR we use the same ones.
@@
Table 14: Hyperparameters of GCN. All models use $2$ layers, a hidden dimension of $64$ and dropout with $0.5$. Different type of weight decay and early stopping from (Gasteiger et al., 2019) was used, if these provided a better performance then using the standard setting in Chien et al. (2021). The same holds for feature normalization, which was used by default in Chien et al. (2021) for GPRGNN.
@@
Table 15: Hyperparameters of GCN for the larger heterophilic datasets. All models use a hidden dimension of $64$, batch norm, no weight decay and the early stopping from Chien et al. (2021).
@@
Dataset DIGL DIGL +JDR #iter FoSR
@@ A.8 Hardware Specifications
Experiments on cSBM, Cora, Citeseer and Photo were conducted on an internal cluster with Nvidia Tesla V100 GPUs with $32$GB of VRAM. The experiments on the remaining datasets (PubMed, Computers, Chameleon, Squirrel, Actor, Cornell and Texas) were performed using Nvidia A100 GPUs with $40$GB or $80$GB of VRAM. The larger VRAM is only necessary for GNN +JDR on PubMed and the larger heterophilic datasets from Lim et al. (2021); Platonov et al. (2023), because they have larger numbers of nodes in the graph (and we choose the top-$64$ edges per node after rewiring). Note that this could be reduced by sacrificing only a little bit of performance as shown in A.6. One experiment of training and testing on $100$ random splits typically takes about $5\min$. For the standard benchmark graphs, the longest experiments with GPRGNN +JDR and a different early stopping condition take about $40\min$. The experiments on the large Twitch-gamers dataset take around $60\min$ (for $10$ splits), but similar to DIGL they require a lot of standard memory (around $500$GB) while performing the decompositions.
A.9 Random Matrix Theory Insight
Following the derivation from Shi et al. (2024), we show empirically how the denoising can reduce the empirical risk for a one-layer GCN without non-linearity. When the number of nodes $N$ goes to infinity and the average node degree satisfies some assumptions, we can apply the Gaussian adjacency equivalence conjecture. This allows us to replace the binary adjacency in the cSBM with a spiked non-symmetric Gaussian random matrix without changing the training and test loss in the limit. The equivalent adjacency reads
@@ A.9 Random Matrix Theory Insight
Figure 16: Experimental results on a non-symmetric gaussian cSBM with $N=1000$ and $\gamma=2$ with denoising of $\bm{A}$. We plot the MSE for different $\mu^{2}$ and $\lambda=0.0$ in 16(a) and $\lambda=1.0$ in 16(b). Each data point is averaged over 10 independent trials and the standard deviation is indicated by the light shaded area.
@@
Figure 17: Experimental results on a non-symmetric gaussian cSBM with $N=1000$ and $\gamma=2$ with denoising of $\bm{X}$. We plot the MSE for different $\lambda$ and $\mu=0.0$ in 17(a) and $\mu=1.0$ in 17(b). Each data point is averaged over 10 independent trials and the standard deviation is indicated by the light shaded area.

### v3

@@ A.5.4 Combining JDR and DIGL
We compare our method to DIGL [Gasteiger et al., 2019] in the main text. We use the personalized PageRank (PPR) diffusion kernel and the same top-$64$ values sparsening method as in JDR in all experiments. Figure 7 shows the additional results for DIGL on the synthetic datasets from the cSBM. Table 12 shows the results on the real-world homophilic datasets in the sparse splitting and Table 13 on the heterophilic datasets in the dense splitting. Here, in addition to the individual results for JDR and DIGL, the results for a combination of the two methods are also shown. For this purpose, the graph was first denoised with JDR and then diffused with DIGL. To do this, we fixed the hyperparameters of JDR and then tuned the parameter $\alpha$ of DIGL. We think this is interesting as both methods enhance the graph in different ways and thus should be combinable. In principle, this should also be possible for a combination of JDR with BORF or FoSR, but so far we have not been able to get this to work.
@@
A.5.5 Alignment
Here, we give a more detailed view on how much JDR actually increases alignment on cSBM and real-world datasets compared to the baseline methods. For cSBM, we can see in Figure 8 that DIGL only increases alignment in the homophilic regime. In the heterophilic regime it clearly decreases alignment. We expect this because it promotes connections among nodes at short diffusion distance. Also the random teleport probability found on these datasets is $1.0$, which results in a random uniformly connected graph. Similar to the results on the real-world datasets Cornell and Texas from the main text, we can see this in the classification performance in Figure 7. In the heterophilic regime, the performance of DIGL matches exactly the MLP, while in the homophilic regime, we can see some performance increases. With GPRGNN, DIGL is not really able to improve performance at all (except for $\phi=0$). In Figure 9, we can see that JDR increases alignment across all settings and more strongly than the baseline methods on the real-world graphs (except for Citeseer). DIGL also increases alignment on many homophilic graphs and on heteophilic Texas. We would like to note that when only rewiring a graph, increasing the alignment might not always be the best thing to do: If the graph is very good, it would be the features that should be made more aligned to the graph. But only JDR is able to do this, as it also denoises the features as well. It is also interesting that DIGL decreases the alignment for Cora and PubMed, but still achieves a good performance. This indicates that DIGL in this case improves the graph in a different way than JDR. So here, it should be possible to combine both methods to achieve even better performance. And indeed, A.5.4 shows that this is exactly the case for Cora, PubMed and Photo on GCN. FoSR and BORF do not visibly change the alignment, since their modifications to the graph are usually too small.
@@ A.5.5 Alignment
Figure 9: Alignment of the leading eigenspaces of graphs from homophilic (9(a)) and heterophilic (9(b)) real-world datasets. We compare the original graph (None) to the output of DIGL, BORF and JDR with the hyperparameters found on GCN. JDR increases the alignment in all settings and achieves the maximum alignment among rewiring algorithms in all settings except on the Citeseer dataset.
A.5.6 Spectral Clustering
@@
Figure 10: Separate results for using spectral clustering on a rewriting only $\bm{A}$ (10(a)) and denoising only $\bm{X}$ (10(b)) compared to full JDR. Note that for $\phi\in\{0.5,0.625,0.875\}$ we had to use additional graphs generated using cSBM with an average node degree of $d=10$ for spectral clustering of $\bm{A}$ to work in general and for $\phi=0.875$ also for JDR. The reason for this is that the graph is very sparse so it is not necessarily connected such that there is no guarantee that spectral clustering works. However a larger node degree does not improve the performance of spectral clustering in general, while it may for GNNs.
@@
Figure 11: Separate results for rewiring only $\bm{A}$ and denoising only $\bm{X}$ compared to full JDR. Results for (11(a)) GCN and (11(b)) GPRGNN. The error bars indicate the $95\%$ confidence interval.
Figure 12: ESNR for graphs from the cSBM with different $\phi$. We compare JDR to DIGL and no rewiring.
A.5.7 Evaluating graph denoising via ESNR
As described in the main text, the Graph Propensity Score (GPS) algorithm [Dong and Kluger, 2023], together with the edge signal-to-noise ratio (ESNR) metric introduced in the same paper, are related to our work. Similar to the GPS algorithm, which uses the ESNR to denoise the graph, JDR denoises the graph but also the features. These two strategies are different, but they are both based on the cSBM. The main drawback of GPS is that it does not consider the possibility to also denoise the features using the graph and it is further limited to low-SNR graphs and to GCN as downstream model. We are interested in the the ESNR metric as it quantifies the noise in the graph.
Table 16: ESNR results of the original and rewired homophilic real-world graphs using JDR and the baseline rewiring methods. Largest value in bold.
Method Cora CiteSeer PubMed Computers Photo
None $0.2964$ $0.1976$ $0.2701$ $0.6502$ $0.7145$
DIGL $\bm{0.6888}$ $\bm{0.5989}$ $\bm{0.5490}$ $\bm{0.7368}$ $\bm{0.7407}$
FoSR $0.2965$ $0.1988$ $0.2700$ $0.6500$ $0.7145$
BORF $0.2958$ $0.1948$ $0.2700$ $0.6502$ $0.7144$
JDR $0.6160$ $0.5754$ $0.5291$ $0.7069$ $0.7155$
Table 17: ESNR results of the original and rewired heterophilic real-world graphs using JDR and the baseline rewiring methods. Largest value in bold.
Method Chameleon Squirrel Actor Texas Cornell
None $0.5199$ $\bm{0.4680}$ $0.0546$ $0.0585$ $0.0388$
DIGL $0.5082$ $0.3884$ $0.0244$ $0.1442$ $0.1195$
FoSR $\bm{0.5199}$ $0.4679$ $0.0546$ $0.0620$ $0.0389$
BORF $0.5199$ $0.0545$ $0.0593$ $0.0423$
JDR $0.4473$ $0.2156$ $\bm{0.1703}$ $\bm{0.2729}$ $\bm{0.2828}$
In Figure 12, Table 16 and Table 17 we compare the achieved ESNR values of JDR with the baselines and the original graphs. The results for cSBM in Figure 12 show that JDR is able to denoise the graph in all cases and DIGL in the strongly homophilic regime. However, comparing the ESNR curve to the actual GCN performance in Figure 7, there is no clear connection apart from the general trend. This was claimed to be more visible in the paper by Dong and Kluger [2023], especially for cSBM on which it is based on. The results on homophilic real-world datasets in Table 16 follow this trend. Indeed both JDR and DIGL are able to decrease the noise in the graph, but DIGL does so best on all five datasets, while the GCN results show a benefit of JDR over DIGL on four out of five datasets (see Table 2). For the heterophilic real-world datasets the results are even more inconsistent. JDR is able to decrease the graph noise on three datasets, but two of them not being the ones where GCN performs best (see Table 3). Also on Chameleon on Squirrel, no method is able to improve the ESNR. We suspect that this behavior occurs due to the role of features which is not captured by the ESNR. Checking the results in the paper by Dong and Kluger [2023], there are cases where the ESNR is not sensitive, e.g. on the Chameleon dataset, which suggests that in such cases, denoising the features is more beneficial than focusing only on the graph structure. In fact, this is in line with our findings in the ablations in Figure 13 and Figure 14, where the denoising of the features improves more than the denoising of the graph on the Chameleon dataset. Overall, the ESNR is generally able to quantify the denoising of the graphs in most cases, but a more direct connection to GCN performance requires further research.
Finally, there are significant differences between the GPS approach by Dong and Kluger [2023] and our work, most importantly in that we consider both graph and feature information. Moreover, JDR improves performance of GNNs on real-world graph datasets where GPS does not provide any improvements (homophilic datasets). Ablations on these datasets (again Figure 13 and Figure 14) indicate that JDR still achieves the most improvement by denoising the graph rather than the features in a way that does not seem to be captured by the ESNR. We think that this is because ESNR does not consider the SNR of the features. In the experiments of Dong and Kluger [2023] with edge dropout on real-world data, the performance barely decreases for datasets like PubMed. The reason for this is that the features already contain a lot of information and making the graph noisier via edge dropout does not spoil this. This is also related to the discussion of GNN vs. MLP (see A.5.3), where an MLP on PubMed already performs reasonably well (see e.g. Table 10.
@@ A.7 Hyperparameters
In this section we list all the hyperparameters used for the experiments to ensure the reproducibility of the results. They are also included in the code. In all experiments we use the Adam optimizer and the standard early stopping after 200 epochs from [Chien et al., 2021]. Whenever we use a GCN, it uses two layers, a hidden dimension of $64$ and dropout with $0.5$. Whenever we use GPRGNN, we use a polynomial filter of order $10$ (corresponding to $10$ hops) and a hidden dimension of $64$. For JDR, we always keep the $64$ largest entries of the rewired adjacency matrix $\tilde{\bm{A}}$ per node. We justify this choice by the ablation in Figure 15.
cSBM. For synthetic data from the cSBM, we generally follow the hyperparameters from [Chien et al., 2021]. GCN uses a learning rate of $0.01$ and weight decay with $\lambda=0.0005$. GPRGNN also uses a $\lambda=0.0005$ and both use ReLU non-linearity. On homophilic graphs ($\phi\geq 0$), GPRGNN uses a learning rate of $0.01$, a weight initialization $\alpha=0.1$ and dropout with $0.5$. For heterophilic graphs, it uses a learning rate of $0.05$, $\alpha=1.0$ and dropout $0.7$. The hyperparameters for JDR on the cSBM are shown in Table 22. We only tuned them using GCN as a downstream model, so for GPRGNN +JDR we use the same ones.
@@
Table 19: Hyperparameters of GCN. All models use $2$ layers, a hidden dimension of $64$ and dropout with $0.5$. Different type of weight decay and early stopping from [Gasteiger et al., 2019] was used, if these provided a better performance then using the standard setting in Chien et al. [2021]. The same holds for feature normalization, which was used by default in Chien et al. [2021] for GPRGNN.
@@
Table 20: Hyperparameters of GCN for the larger heterophilic datasets. All models use a hidden dimension of $64$, batch norm, no weight decay and the early stopping from Chien et al. [2021].
@@
Dataset $\alpha$ DIGL $\alpha$ DIGL +JDR #iter FoSR
@@ A.8 Hardware Specifications
Experiments on cSBM, Cora, Citeseer and Photo were conducted on an internal cluster with Nvidia Tesla V100 GPUs with $32$GB of VRAM. The experiments on the remaining datasets (PubMed, Computers, Chameleon, Squirrel, Actor, Cornell and Texas) were performed using Nvidia A100 GPUs with $40$GB or $80$GB of VRAM. The larger VRAM is only necessary for GNN +JDR on PubMed and the larger heterophilic datasets from Lim et al. [2021], Platonov et al. [2023], because they have larger numbers of nodes in the graph (and we choose the top-$64$ edges per node after rewiring). Note that this could be reduced by sacrificing only a little bit of performance as shown in A.6. One experiment of training and testing on $100$ random splits typically takes about $5\min$. For the standard benchmark graphs, the longest experiments with GPRGNN +JDR and a different early stopping condition take about $40\min$. The experiments on the large Twitch-gamers dataset take around $60\min$ (for $10$ splits), but similar to DIGL they require a lot of standard memory (around $500$GB) while performing the decompositions.
A.9 Insights from Random Matrix Theory for one-layer GCNs
Following the derivation from Shi et al. [2024], we show empirically how denoising can reduce the empirical risk for a one-layer GCN without non-linearity. When the number of nodes $N$ goes to infinity and the average node degree satisfies some assumptions, we can apply the Gaussian adjacency equivalence conjecture. This allows us to replace the binary adjacency in the cSBM with a spiked non-symmetric Gaussian random matrix without changing the training and test loss in the limit. The equivalent adjacency reads
@@ A.9 Insights from Random Matrix Theory for one-layer GCNs
Figure 18: Experimental results on a non-symmetric gaussian cSBM with $N=1000$ and $\gamma=2$ with denoising of $\bm{A}$. We plot the MSE for different $\mu^{2}$ and $\lambda=0.0$ in 18(a) and $\lambda=1.0$ in 18(b). Each data point is averaged over 10 independent trials and the standard deviation is indicated by the light shaded area.
@@
Figure 19: Experimental results on a non-symmetric gaussian cSBM with $N=1000$ and $\gamma=2$ with denoising of $\bm{X}$. We plot the MSE for different $\lambda$ and $\mu=0.0$ in 19(a) and $\mu=1.0$ in 19(b). Each data point is averaged over 10 independent trials and the standard deviation is indicated by the light shaded area.